A dog of mass 5Kg is standing on a flat boat so that he is 10m from the shore.He walks 4m on the boat towards the shore and then halts. If the mass of the boat is 20Kg, find the distance of the dog from the shore when it halts________

Grade:11

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

let the x1 is the distance of center of mass of boat from the shore,and Xc.m is the distance of  center of mass of the system from the shore

Xc.m =(5*10 +20*x1)/(20+5)

          =2 +(4/5)  x1

now let boat move x distance away from the shore when dog move 4m towards the shore

so new distance of boat from shore=x1 +x

new distance of dog from shore=10-4 +x=6+x

now

 Xc.m ={5*(6+x) +20*(x1+x)}/(20+5)

          =(30+25x +20x1)/25

          =6/5 +x +4/5x1

 

equate both Xc.m

6/5 +x +4/5x1 =2 +(4/5)  x1

 x=2-6/5

  =4/5

so distance =6+4/5

                 =34/5


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

Raunaq Mehta
7 years ago
let the x1 is the distance of center of mass of boat from the shore,and Xc.m is the distance of center of mass of the system from the shoreXc.m =(5*10 +20*x1)/(20+5) =2 +(4/5) x1now let boat move x distance away from the shore when dog move 4m towards the shoreso new distance of boat from shore=x1 +xnew distance of dog from shore=10-4 +x=6+xnow Xc.m ={5*(6+x) +20*(x1+x)}/(20+5) =(30+25x +20x1)/25 =6/5 +x +4/5x1 equate both Xc.m6/5 +x +4/5x1 =2 +(4/5) x1 x=2-6/5 =4/5so distance =6+4/5 =34/5